a steam turbine receives steam at 1379 kpa gauge and exhausts it to a condenser where the vacuum is indicated as 635 mm of hg by a gauge. for a barometric pressure of 760 mm of Hg. calculate a) the absolute pressure of the turbine inlet, kpa b) the absolute pressure corresponding to the vacuum gauge reading, kpa and mm of hg
1
Expert's answer
2020-02-27T10:04:18-0500

As per the given question,

gauge pressure =1379kpa

Vaccume Pressure =635mm of hg

Barometric pressure=760 mm of hg

Patm=101.3kpa

As we know,



a)

"P_{inlet}=P_{atm}+P_{gauge}"

"P_{inlet}=(101.3+1379)Kpa"

"=1480.3Kpa"

b)

"P_{absolute}=P_{barometric}+P_{vaccume}"

"=760-635=125mm" of hg

Now converting mm hg pressure to Kpa

1 cm Hg = 1333.2239 pascal

"So, 1mm hg=\dfrac{1333.2239}{10}=133.33pa"

So, 125mm of hg="125\times133.33pa=16666.25pa=16.67Kpa"